Evalute: 1) sin 30 + tan 45 - Cosec 60/ sec 30 + cos 60 + cot 45

Evalute: 
1) sin 30 + tan 45 - Cosec 60/ sec 30 + cos 60 + cot 45

Grade:10

1 Answers

Harshit Singh
3 years ago
Dear Student

sin 30 = 1/2
tan 45 = 1
cosec 60 = 2/√3
sec30 = 2/√3
cos60 = 1/2
cot45=1
Putting all these values
sin 30 + tan 45 - Cosec 60/ sec 30 + cos 60 + cot 45= ½+1 –1 + ½ +1
= 2

Thanks